The Qd represents the amount consumed of good X. Px is the price of good X, Py, is the price of good Y, M is the level of income, and A is the level of expenditure in the advertisement. Assuming good X sells for P40 per unit and Y sells at P30/unit, the firm uses 4,000 units of advertising, and the target buyer average income is P10,000. How much of good X will consumers buy? Is good Y a substitute to X or complements? Is good X a normal or inferior good?
1
Expert's answer
2020-09-14T10:47:27-0400

Solution:

1.). The amount of good X that consumers will buy:

"Q_{d} =a-b(P)"


Where: a = 4000

"b=\frac{Px}{Py} =\frac{30}{40} = 0.75"


"Q_{d}=4000-0.75(40)"

"= 4000-30"

"=3970\; units"

2.). Good Y is a substitute of good X, since it has a positive cross elasticity of demand, that is the goods demand is increased when the price is increased.


3.). Good X is a normal good since it has a positive income elasticity, that is the percentage change in income leads to a percentage change in quantity demanded.